# revtex4-2Repair the float

### v3

@@ 4 PDW Bogoliubov-de Gennes Hamiltonian with Topological defects
In this section we describe the Bogoliubov-de Gennes Hamiltonian on the square lattice with a cuprate electronic structure with the configurations of the PDW order parameter in the background of the topological defects introduced in sections 2 and 3. We will focus on the effects on the electronic states.
@@
$\Delta_{i}(\mathbf{r},\mathbf{r}^{\prime})=\bar{\Delta}\,F(\mathbf{r},\mathbf{r}^{\prime})f_{i}(\mathbf{r})$ (4.1)
@@
Here $\bar{\Delta}$ is the amplitude of the SC gap given in Eq.(3.6), $F(\mathbf{r},\mathbf{r}^{\prime})$ is the SC form factor and $f_{i}(\mathbf{r})$ are the profiles and winding-numbers of the four configurations of the PDW order parameters listed above. On a square lattice the form factor $F(\bm{r},\bm{r}^{\prime})=1$ for an s-wave SC state. In a d-wave SC state, which is our focus, the form factor is $F(\bm{r},\bm{r}^{\prime})=1(-1)$ for a bond $(\bm{r},\bm{r}^{\prime})$ on the x axis (y axis) of the square lattice, and changes sign under a $\pi/2$ rotation. Using the results of section 3 the explicit forms of the functions $f_{i}(\mathbf{r})$’s are
$f_{1}(\mathbf{r})=$ $\cos(\mathbf{Q}\cdot\mathbf{r})$ (4.2)
$f_{2}(\mathbf{r})=$ $\frac{1}{2}\Big{(}\sin\big{(}\alpha(\mathbf{r})\big{)}e^{i\mathbf{Q}\cdot\mathbf{r}+i\varphi(\mathbf{r})}+\cos\big{(}\alpha(\mathbf{r})\big{)}e^{-i\mathbf{Q}\cdot\mathbf{r}}\Big{)}$
$f_{3}(\mathbf{r})=$ $f(r/r_{0})\cos(\mathbf{Q}\cdot\mathbf{r})e^{i\varphi(\mathbf{r})}$
$f_{4}(\mathbf{r})=$ $f(r/r_{0})\cos(\mathbf{Q}\cdot\mathbf{r}+\varphi(\mathbf{r}))$
@@
Figure 3: (Online Color) Plots of the real parts of the daughter CDWs $\rho_{2\mathbf{Q}}(\mathbf{r})$ given by equation (2.6) defined in terms of the functions given in equation (4.2). Here the three topological defects have a vortex radius of $r_{0}=16a_{0}$ being: (a) the half-vortex, (b) the Abrikosov vortex and (c) the double dislocation. The dotted lines shown in light grey are guides to count the CDW peaks to find the associated Burgers vectors, being the difference of the top line and the bottom, As expected, in (a) for the half-vortex (which has a single dislocation) they skip by one, in (b) for the Abrikosov vortex they do not skip, and in (c) for the double dislocation they skip by two. In the second row, plots (d)-(f), we include the corresponding $\text{arg}\big{(}\rho_{2\mathbf{Q}}\big{)}$ for (d) the half-vortex, (e) the Abrikosov vortex and (f) the double dislocation. The jumps in phase seen in these figures are $\pi/2$ and they sum up to the expected dislocation charge associated with a given defect. The color bar provided in (g) pertains to all plots. For (a)-(c) it corresponds to the scale of the defect (that is we normalized these plots), and for (d)-(f) it represents units of $2\pi$.
@@
Figure 4: (Online Color) The LDOS profiles of a PDW in the presence of the three topological defects given in Eq. (4.2). Using the states of the BdG Hamiltonian (Eq.(4.3)) the LDOS evaluated at $\omega=.25\;\bar{\Delta}$ for (a) the half-vortex, (b) the Abrikosov vortex and (c) the double dislocation. In (d) we provide a normalized color bar for each of the plots. Note that each specific defect is rescaled according to their own maximum value, so that the vortex core shows up clearer.
@@
Here $\alpha(\mathbf{r})$ is the angle we used to parameterize the NLSM in Eq.(3.9), whose numerical solution was found in section 3. The vortex profile function $f(r/r_{0})$ is defined in Eq.(3.17). Finally, the complex phase, $\varphi(\mathbf{r})$ is the azimuthal angle on the plane and it winds by $2\pi$ in all the expressions it appears. In figures 3a-3c we show the profile of the composite order parameters $\rho_{2\mathbf{Q}}(\mathbf{r})$ (that is we take the real part of this expression) in the presence of the three topological defects of our PDW order. The Burgers vector associated with a given charge dislocation can be found by simply counting the difference in the CDW peaks found above and below the vortex cores. The dotted lines are guides used to indicate where to do the counting. The profiles of the associated CDW order in the presence of the defects are shown in section 4.2. We also include $\text{arg}(\rho_{2\mathbf{Q}})$ in figures 3d-3f. Notice the four jumps in phase by $\pi/2$ for the half-vortex (figure 3d) and the eight jumps for the double dislocation (figure 3f), while there are none for the full vortex (figure 3e), reflecting the expected amount of dislocation charge present in each defect.
@@ 4.1 Hamiltonian and Observables
for each configuration of the SC amplitudes $\Delta_{i}(\bm{r},\bm{r}^{\prime})$ (see equations 4.1 and 4.2). The normal state band structure we will be using is parameterized with values of hopping amplitudes of a tight-binding model on the square lattice chosen to best fit Angle-Resolved-Photoemission-spectroscopy (ARPES) experiments in the high temperature superconductors La2-xBaxCuO4 and Bi2Sr2CaCu2O8+δ [61, 62]. The explicit parameters used (in units of eV) are: $t=0.25,\;t^{{}^{\prime}}=-0.031863,\;t^{{}^{\prime\prime}}=0.016487,\;t^{{}^{\prime\prime\prime}}=0.0076112$, and $\mu=-0.16235$. We also take the superconducting amplitude $\bar{\Delta}=60$ meV. In all cases we assumed that the superconducting order parameter $\Delta(\bm{r},\bm{r}^{\prime})$ is a unidirectional PDW along the $x$ direction with period $8$ lattice spacings, with a wave vector $\bm{Q}=(\pi/4,0)$ (in units with $a=1$).
Since the Fermi surface of the cuprates is not spherically symmetric, the PDW states along the nodal and anti-nodal directions have different features. The same applies for a putative FF state. Below we will show the Bogoliubov spectrum in the core of the half-vortex resembles that of a pristine FF state trapped inside. For a wave vector oriented along the anti-nodal direction results in a fully gapped FF state, whereas for a state oriented in the nodal direction the resulting spectrum has nodes. In the situation of interest the FF state in the core of the half-vortex of the PDW is gapped.
The details of the diagonalization procedure can be found in appendices A and B. In short, we define a Nambu spinor, $\psi_{+}^{T}=[\mathbf{c}_{\uparrow},\mathbf{c}^{\dagger}_{\downarrow}]$, which helps us perform the exact diagonalization (see appendix B). These define our quasi-particle operators, $\hat{b}_{l}$ and $\hat{b}_{l}^{\dagger}$, which annihilate the BCS ground state and create single particle excitations with energy $E_{l}$, respectively. [^1] Here we take as the zero of energy the ground state value, $E_{G}$.
As in the case of a uniform superconductor, the excited states are an admixture of electrons and holes. We find our electron creation/annihilation operators are related to linear combinations of our quasi-particle operators: $\hat{c}_{i\sigma}=v^{*}_{il}\hat{b}_{l}^{\dagger}+\sigma u_{il}\hat{b}_{l}$. Here repeated indices are summed over, and the coefficients are the real space coherence factors.
Figure 5: (Online Color) Static charge density profiles obtained by integrating the LDOS out to a voltage of $1.25\,\bar{\Delta}$ in the presence of the three topological defects given in Eq.(4.2). Using the states of the BdG Hamiltonian (Eq.(4.3)) the LDOS for a PDW defects for (a) the half-vortex, (b) the Abrikosov vortex and (c) the double dislocation. We take a scale normalized to the specific defect for each of these plots.
@@
$=$ $\frac{1}{\pi}\sum_{E_{l}\geq 0}\Bigg{(}\frac{\epsilon}{(\omega-E_{l})^{2}+\epsilon^{2}}|u_{il}|^{2}+\frac{\epsilon}{(\omega+E_{l})^{2}+\epsilon^{2}}|v_{il}|^{2}\Bigg{)}$ (4.4)
@@ 4.2 Electronic Structure of the PDW Topological defects
In this subsection we analyze our numerical results for the LDOS, computed using Eq. (4.4), for the configurations of the PDW order parameter with the three topological defects defined in Eq. (4.2). The intertwining of the PDW defects with the induced CDW order will be discussed in detail, as well as the structure of the charge distribution induced by these defects. The main focus will be on experimental signatures associated with the CDW pattern induced by the half-vortex and the double dislocation. The superconducting properties of the PDW half-vortex will be discussed in subsection 4.4.
@@
In figure 3 we plot the profiles of (i.e. we take the real part of) the resulting CDW order parameter $\rho_{2\bm{Q}}$ near the three topological defects given in equation (4.2) using the definition of equation (2.6). We note that in our numerics we take the form factor $F(\mathbf{r},\mathbf{r}^{\prime})$ to be defect-free d-wave. Since the PDW order breaks the point group symmetry of the lattice, the form factor associated with the unidirectional PDW phase should be an admixture of s-wave and d-wave [16]. However, as was discussed in [64], there are robust features which are essentially the same for both form factors. In appendix F figure 14 we present the spectral functions for an order parameters with an s-wave form factor, but our primary focus will be on d-wave SC.
@@
In figure 4 we show the changes in the LDOS of a the PDW state with the three topological defects whose CDW order parameter $\rho_{2\bm{Q}}$ near the defects are shown in figure 3. The LDOS of these defects are shown in figures 4a-4c for a probing voltage of $0.25\Delta_{0}$. These were obtained by computing numerically the tunneling density of states of the electronic states obtained from the BdG equations for the three defects. These patterns exhibit a sinusoidal PDW oscillatory component of four lattice spacings, as expected for a CDW with ordering wave vector $\bm{Q}_{cdw}=2\bm{Q}$ (see Eq.(2.6)), superposed with various effects arising from the changes induced by the topological defects on the eigenstates of the BdG equation.
@@
The charge density profiles associated with each of these defects reveals some of the most salient signatures of the PDW order. First and foremost, figure 4a shows that the half-vortex which can indeed be thought of as a dislocation in the CDW order parameter $\rho_{2\bm{Q}}(\bm{r})$ pinned to a half-SC-flux-quanta. The predicted forms of the other two topological defects have been discussed in the literature  [45, 44, 2]. The double dislocation is shown in figure 4c. As was the case for figure 3, the Burgers vector associated with these charge dislocations can be found by simply counting the difference in the CDW peaks found above and below the vortex cores. The full-vortex has no dislocation charge (figure 4b). Notice however the phase of background density wave pattern of the full-vortex is shifted by $\pi$ relative to the pattern of the other two defects, which can easily be seen at $x/a_{0}=0$.
@@
In figure 5 we plot the integrated LDOS (i.e. the static charge density) for all three defects to a voltage of $1.25\,\bar{\Delta}$, well above the PDW SC gap. A comparison of these plots with figure 3 shows, as expected, that the integrated LDOS yields the CDW pattern (for details of this approach see Ref.[56]). As expected, in both figures the CDW order parameter is suppressed in the core of the defects where one or both components of the PDW order parameters $\Delta_{\pm\bm{Q}}$ are suppressed.
@@
Next we notice the additional patterns seen within the core of the defects in figures 4b-4c, the Abrikosov vortex and the double dislocation. Both PDW order parameters $\Delta_{\pm\bm{Q}}$ vanish in the core of the double dislocation and of the Abrikosov vortex; hence, the additional electronic structure residing in their cores, revealed by the LDOS, is due to quasi-particle states. Although the quasi-particle states are responsible for the additional LDOS they are not bound to the core of these two types of topological defects. The PDW has pockets of quasi-particle and quasi-holes in momentum space. This interpretation is confirmed by a computation of the Fourier transforms in momentum space of the LDOS at different energies for the Abrikosov vortex. Figure 13 in appendix D shows the quasi-particle spectrum in the presence of the vortex which confirms that these are propagating states and are not bound to the core of the defect. Thus this structure in the LDOS has to be interpreted as due to quasi-particle interference (QPI) at the defects.
@@
On the other hand, the half-vortex in figure 4a does not posses the QPI patterns seen for the double dislocation and the Abrikosov vortex. This is because in the case of the half-vortex one component of the PDW is always nonzero which results in a gap for the states within the half-vortex core where the PDW becomes effectively a fully gapped FF-state. In contrast, in the cases of the Abrikosov vortex and of the double dislocation both components of the PDW order parameter, $\Delta_{\pm\bm{Q}}$, vanish at the core and the BdG states become gapless at there. Other details associated with LDOS of our defects can be found in appendix D where a zero bias probing voltages is considered.

### v4

@@ 4 PDW Bogoliubov-de Gennes Hamiltonian with Topological defects
In this section we describe the Bogoliubov-de Gennes Hamiltonian on the square lattice with a cuprate electronic structure with the configurations of the PDW order parameter in the background of the topological defects introduced in Secs. 2 and 3. We will focus on the effects on the electronic states.
@@
$\Delta_{i}(\mathbf{r},\mathbf{r}^{\prime})=\bar{\Delta}\,F(\mathbf{r},\mathbf{r}^{\prime})f_{i}(\mathbf{r}).$ (4.1)
@@
Here $\bar{\Delta}$ is the amplitude of the SC gap given in Eq. (3.6), $F(\mathbf{r},\mathbf{r}^{\prime})$ is the SC form factor, and $f_{i}(\mathbf{r})$ are the profiles and winding numbers of the four configurations of the PDW order parameters listed above. On a square lattice the form factor $F(\bm{r},\bm{r}^{\prime})=1$ for an $s$-wave SC state. In a $d$-wave SC state, which is our focus, the form factor is $F(\bm{r},\bm{r}^{\prime})=1(-1)$ for a bond $(\bm{r},\bm{r}^{\prime})$ on the $x$ axis ($y$ axis) of the square lattice, and changes sign under a $\pi/2$ rotation. Using the results of section 3 the explicit forms of the functions $f_{i}(\mathbf{r})$’s are
$f_{1}(\mathbf{r})=$ $\cos(\mathbf{Q}\cdot\mathbf{r}),$ (4.2)
$f_{2}(\mathbf{r})=$ $\frac{1}{2}\Big{(}\sin\big{(}\alpha(\mathbf{r})\big{)}e^{i\mathbf{Q}\cdot\mathbf{r}+i\varphi(\mathbf{r})}+\cos\big{(}\alpha(\mathbf{r})\big{)}e^{-i\mathbf{Q}\cdot\mathbf{r}}\Big{)},$
$f_{3}(\mathbf{r})=$ $f(r/r_{0})\cos(\mathbf{Q}\cdot\mathbf{r})e^{i\varphi(\mathbf{r})},$
$f_{4}(\mathbf{r})=$ $f(r/r_{0})\cos[\mathbf{Q}\cdot\mathbf{r}+\varphi(\mathbf{r})].$
@@
Figure 3: (Online Color) Plots of the real parts of the daughter CDWs $\rho_{2\mathbf{Q}}(\mathbf{r})$ given by Eq.  (2.6) defined in terms of the functions given in Eq.  (4.2). Here the three topological defects have a vortex radius of $r_{0}=16a_{0}$ being: (a) the half-vortex, (b) the Abrikosov vortex, and (c) the double dislocation. The dotted lines shown in light gray are guides to count the CDW peaks to find the associated Burgers vectors, being the difference of the top line and the bottom, As expected, in (a) for the half-vortex (which has a single dislocation) they skip by one, in (b) for the Abrikosov vortex they do not skip, and in (c) for the double dislocation they skip by two. In the second row, plots (d)-(f), we include the corresponding $\text{arg}\big{(}\rho_{2\mathbf{Q}}\big{)}$ for (d) the half-vortex, (e) the Abrikosov vortex and (f) the double dislocation. The jumps in phase seen in these panels are $\pi/2$, and they sum up to the expected dislocation charge associated with a given defect. The color bar provided in (g) pertains to all plots. For (a)-(c) it corresponds to the scale of the defect (that is, we normalized these plots), and for (d)-(f) it represents units of $2\pi$.
@@
Figure 4: (Online Color) The LDOS profiles of a PDW in the presence of the three topological defects given in Eq. (4.2). Using the states of the BdG Hamiltonian [Eq.(4.3)] the LDOS evaluated at $\omega=0.25\;\bar{\Delta}$ for (a) the half-vortex, (b) the Abrikosov vortex, and (c) the double dislocation. In (d) we provide a normalized color bar for each of the plots. Note that each specific defect is rescaled according to their own maximum value, so that the vortex core shows up clearer.
@@
Here $\alpha(\mathbf{r})$ is the angle we used to parametrize the NLSM in Eq.(3.9), whose numerical solution was found in Sec. 3. The vortex profile function $f(r/r_{0})$ is defined in Eq. (3.17). Finally, the complex phase, $\varphi(\mathbf{r})$, is the azimuthal angle on the plane, and it winds by $2\pi$ in all the expressions in which it appears. In Figs. 3a-3c we show the profile of the composite order parameters $\rho_{2\mathbf{Q}}(\mathbf{r})$ (that is, we take the real part of this expression) in the presence of the three topological defects of our PDW order. The Burgers vector associated with a given charge dislocation can be found by simply counting the difference in the CDW peaks found above and below the vortex cores. The dotted lines are guides used to indicate where to do the counting. The profiles of the associated CDW order in the presence of the defects are shown in Sec. 4.2. We also include $\text{arg}(\rho_{2\mathbf{Q}})$ in Figs. 3d-3f. Notice the four jumps in phase by $\pi/2$ for the half-vortex [Fig. 3d] and the eight jumps for the double dislocation [Fig. 3f], while there are none for the full vortex [Fig. 3e], reflecting the expected amount of dislocation charge present in each defect.
@@ 4.1 Hamiltonian and Observables
for each configuration of the SC amplitudes $\Delta_{i}(\bm{r},\bm{r}^{\prime})$ [see Eqs. (4.1) and (4.2)]. The normal state band structure we will be using is parameterized with values of hopping amplitudes of a tight-binding model on the square lattice chosen to best fit Angle-Resolved Photoemission Spectroscopy (ARPES) experiments in the high-temperature superconductors La2-xBaxCuO4 and Bi2Sr2CaCu2O8+δ [60, 61]. The explicit parameters used (in units of eV) are: $t=0.25,\;t^{{}^{\prime}}=-0.031863,\;t^{{}^{\prime\prime}}=0.016487,\;t^{{}^{\prime\prime\prime}}=0.0076112$, and $\mu=-0.16235$. We also take the superconducting amplitude $\bar{\Delta}=60$ meV. In all cases we assumed that the superconducting order parameter $\Delta(\bm{r},\bm{r}^{\prime})$ is a unidirectional PDW along the $x$ direction with period eight lattice spacings, with a wave vector $\bm{Q}=(\pi/4,0)$ (in units with $a=1$).
Since the Fermi surface of the cuprates is not spherically symmetric, the PDW states along the nodal and anti-nodal directions have different features. The same applies for a putative FF state. Below we will show the Bogoliubov spectrum in the core of the half-vortex resembles that of a pristine FF state trapped inside. A wave vector oriented along the anti-nodal direction results in a fully gapped FF state, whereas for a state oriented in the nodal direction the resulting spectrum has nodes. In the situation of interest the FF state in the core of the half-vortex of the PDW is gapped.
The details of the diagonalization procedure can be found in Appendixes A and B. In short, we define a Nambu spinor, $\psi_{+}^{T}=[\mathbf{c}_{\uparrow},\mathbf{c}^{\dagger}_{\downarrow}]$, which helps us perform the exact diagonalization (see Appendix B). These define our quasiparticle operators, $\hat{b}_{l}$ and $\hat{b}_{l}^{\dagger}$, which annihilate the BCS ground state and create single particle excitations with energy $E_{l}$, respectively  [^1] Here we take as the zero of energy the ground state value, $E_{G}$..
As in the case of a uniform superconductor, the excited states are an admixture of electrons and holes. We find our electron creation/annihilation operators are related to linear combinations of our quasiparticle operators: $\hat{c}_{i\sigma}=v^{*}_{il}\hat{b}_{l}^{\dagger}+\sigma u_{il}\hat{b}_{l}$. Here repeated indices are summed over, and the coefficients are the real space coherence factors.
Figure 5: (Online Color) Static charge density profiles obtained by integrating the LDOS out to a voltage of $1.25\,\bar{\Delta}$ in the presence of the three topological defects given in Eq.(4.2). Using the states of the BdG Hamiltonian [Eq. ((4.3))] the LDOS for a PDW defects for (a) the half-vortex, (b) the Abrikosov vortex, and (c) the double dislocation. We take a scale normalized to the specific defect for each of these plots.
@@
$=$ $\frac{1}{\pi}\sum_{E_{l}\geq 0}\Bigg{(}\frac{\epsilon}{(\omega-E_{l})^{2}+\epsilon^{2}}|u_{il}|^{2}+\frac{\epsilon}{(\omega+E_{l})^{2}+\epsilon^{2}}|v_{il}|^{2}\Bigg{)},$ (4.4)
@@ 4.2 Electronic Structure of the PDW Topological defects
In this subsection we analyze our numerical results for the LDOS, computed using Eq. (4.4), for the configurations of the PDW order parameter with the three topological defects defined in Eq. (4.2). The intertwining of the PDW defects with the induced CDW order will be discussed in detail, as well as the structure of the charge distribution induced by these defects. The main focus will be on experimental signatures associated with the CDW pattern induced by the half-vortex and the double dislocation. The superconducting properties of the PDW half-vortex will be discussed in Sec. 4.4.
@@
In Fig. 3 we plot the profiles of (i.e., we take the real part of) the resulting CDW order parameter $\rho_{2\bm{Q}}$ near the three topological defects given in Eq.  (4.2) using the definition of Eq.  (2.6). We note that in our numerics we take the form factor $F(\mathbf{r},\mathbf{r}^{\prime})$ to be defect-free $d$-wave. Since the PDW order breaks the point group symmetry of the lattice, the form factor associated with the unidirectional PDW phase should be an admixture of $s$-wave and $d$-wave [16]. However, as was discussed in [63], there are robust features which are essentially the same for both form factors. In Appendix F, Fig. 14, we present the spectral functions for an order parameters with an $s$-wave form factor, but our primary focus will be on $d$-wave SC.
@@
In Fig. 4 we show the changes in the LDOS of a the PDW state with the three topological defects whose CDW order parameter $\rho_{2\bm{Q}}$ near the defects are shown in Fig. 3. The LDOS of these defects are shown in Figs. 4a-4c for a probing voltage of $0.25\Delta_{0}$. These were obtained by computing numerically the tunneling density of states of the electronic states obtained from the BdG equations for the three defects. These patterns exhibit a sinusoidal PDW oscillatory component of four lattice spacings, as expected for a CDW with ordering wave vector $\bm{Q}_{cdw}=2\bm{Q}$ (see Eq.(2.6)), superposed with various effects arising from the changes induced by the topological defects on the eigenstates of the BdG equation.
@@
The charge density profiles associated with each of these defects reveals some of the most salient signatures of the PDW order. First and foremost, Fig. 4a shows that the half-vortex which can indeed be thought of as a dislocation in the CDW order parameter $\rho_{2\bm{Q}}(\bm{r})$ pinned to a half-SC-flux-quanta. The predicted forms of the other two topological defects have been discussed in the literature  [45, 44, 2]. The double dislocation is shown in Fig. 4c. As was the case for Fig. 3, the Burgers vector associated with these charge dislocations can be found by simply counting the difference in the CDW peaks found above and below the vortex cores. The full vortex has no dislocation charge [Fig. 4b]. Notice, however, the phase of background density wave pattern of the full vortex is shifted by $\pi$ relative to the pattern of the other two defects, which can easily be seen at $x/a_{0}=0$.
@@
In Fig. 5 we plot the integrated LDOS (i.e., the static charge density) for all three defects to a voltage of $1.25\,\bar{\Delta}$, well above the PDW SC gap. A comparison of these plots with Fig. 3 shows, as expected, that the integrated LDOS yields the CDW pattern (for details of this approach see Ref. [55]). As expected, in both figures the CDW order parameter is suppressed in the core of the defects where one or both components of the PDW order parameters $\Delta_{\pm\bm{Q}}$ are suppressed.
@@
Next we notice the additional patterns seen within the core of the defects in Figs. 4b and 4c, the Abrikosov vortex and the double dislocation. Both PDW order parameters $\Delta_{\pm\bm{Q}}$ vanish in the core of the double dislocation and of the Abrikosov vortex; hence, the additional electronic structure residing in their cores, revealed by the LDOS, is due to quasiparticle states. Although the quasiparticle states are responsible for the additional LDOS, they are not bound to the core of these two types of topological defects. The PDW has pockets of quasiparticle and quasi-holes in momentum space. This interpretation is confirmed by a computation of the Fourier transforms in momentum space of the LDOS at different energies for the Abrikosov vortex. Figure 13 in Appendix D shows the quasiparticle spectrum in the presence of the vortex, which confirms that these are propagating states and are not bound to the core of the defect. Thus this structure in the LDOS has to be interpreted as due to quasiparticle interference (QPI) at the defects.
@@
On the other hand, the half-vortex in Fig. 4a does not posses the QPI patterns seen for the double dislocation and the Abrikosov vortex. This is because in the case of the half-vortex one component of the PDW is always nonzero, which results in a gap for the states within the half-vortex core where the PDW becomes effectively a fully gapped FF state. In contrast, in the cases of the Abrikosov vortex and of the double dislocation both components of the PDW order parameter, $\Delta_{\pm\bm{Q}}$, vanish at the core, and the BdG states become gapless at there. Other details associated with LDOS of our defects can be found in Appendix D, where a zero bias probing voltages is considered.